OpenStudy (anonymous):

how do I solve 4cos(x+4)=1?

OpenStudy (jdoe0001):

\(\bf 4cos(x+4)=1\implies cos(x+4)=\cfrac{1}{4}\\ \quad \\ \implies cos^{-1}[cos(x+4)]=cos^{-1}\left(\cfrac{1}{4}\right)\)

OpenStudy (jdoe0001):

recall that \(\bf cos^{-1}[cos(\theta)]=\theta\)

OpenStudy (anonymous):

\[Cos(x+4) = \frac{ 1 }{ 4 }\] \[x+4 = Cos ^{-1}(1/4)\] \[x = Cos ^{-1}(1/4) -4\] And you need to use the calculator to find the value of arc cos(1/4)

OpenStudy (anonymous):

okay thank you that helped a lot!